# Generative AI and Firm Productivity:

### v2

@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
(1) (2) (3) (4) (5)
Click Order Sales Conversion Rate Cart Value
Treat 0.00233 0.000600** 0.0104** 0.000554*** 0.0942
(0.00183) (0.000278) (0.00417) (0.000187) (0.0807)
%Change 0.12% 1.08% 2.05% 1.27% 0.81%
Observations 4,772,937 4,772,937 4,772,937 4,772,937 209,371
R-squared 0.055 0.008 0.002 0.008 0.010
@@
    “Click” stands for the number of product clicks into product detail pages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Marketing Push Message.
@@
In Table C5, Column 1 reveals that GenAI-generated content accounts for 40% of the marketing push message in the treatment group. With the use of GenAI-generated messages, there are observable rises in consumer clicks and purchases, with the number of clicks increasing by 3.1%, the number of orders increasing by 2.8%, and the purchase amount growing by 1.6% (Columns 2 to 4). Additionally, the likelihood of a consumer making a purchase increased by 3% in Column 5. The Cart Value is not statistically significant (Column 6). The results highlight the benefits of using GenAI in unlocking the personalization potential of marketing content, particularly in environments where human-generated content alone may be limited by resource constraints.
@@
Is AI Task Click Order Sales Conversion Rate Cart Value
Treat 0.394*** 0.000529** 0.00005* 0.000402 0.000048*** 0.024
@@
%Change 3.1% 2.8% 1.6% 3.0% 0.15%
Observations 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528 11,030
R-squared 0.245 0.000 0.000 0.000 0.000 0.000
@@
    “Is AI Task” is a binary variable which equals one if the message is generated by AI. “Click” stands for the number of clicks on the marketing messages. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@
View Click Sales Conversion Rate Cart Value
@@
    “View” represents the number of times the product advertisement is viewed on Google. “Click” refers to the number of times the product advertisement is clicked on Google. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures users’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per product, conditional on the product being purchased.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@
Appendix D Details on the Heterogeneous Treatment Effects Across Sellers
Search Query Refinement.
@@ Appendix D Details on the Heterogeneous Treatment Effects Across Sellers
Table D1 reports the heterogeneous impact on high- versus low-type of sellers classified based on three metrics: annual past sales, tenure on the platform, and number of sub-accounts. Small sellers with lower transaction volumes (Column 2), shorter operational histories (Column 4), and fewer sub-accounts (Column 6) experience a significant increase in both sales (Panel A) and conversion rates (Panel B) from treated consumers. In contrast, expenditures and conversion rates on big, tenured and scaled sellers show no significant change. Thus, the GenAI-powered search query refinement generates more values among small sellers.
Table D1: Seller HTE for Search Query Refinement
@@
%Change 0.21% 1.69% 1.24% 1.01% 0.88% 1.20%
Observation 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382
@@
    We classify sellers into high and low groups based on three different proxies. Low-group sellers are generally small sellers, as defined by meeting the following pre-experiment criteria: (1) accounting for the bottom 50% cumulative share of total sales when sellers are ranked by annual sales; (2) having operated on the platform for fewer than five years; or (3) maintaining fewer than three sub-accounts for their online store.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Marketing Push Message.
Table D2 exhibited the heterogeneous impact of Marketing Push Message. Similarly, we observe that small sellers experience a significant increase in conversion rates from treated consumers (Columns 2, 4 and 6 in Panel B), while the conversion rates on big sellers show no significant changes between treatment and control groups. Thus, GenAI helps the platform to more effectively leverage content personalization in its marketing activities, particularly benefiting small sellers.
Table D2: Seller HTE for Marketing Push Message
@@
% Change 3.3% 2.7% 0.9% 3.8% -0.1% 4.8%
Observation 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528
@@
    Please refer to Table D1 for detailed notes.
Google Advertising Titles.
Table D3 shows that the heterogeneous treatment effects are statistically insignificant across both high- and low-seller groups, for sales as well as conversion rate.
Table D3: Seller HTE for Google Advertising Titles
@@
% Change -6.5% -0.3% -2.3% -3.8% -4.7% -3.2%
Observations 622,083 621,933 397,339 846,677 140,005 1,104,011
@@
    Please refer to Table D1 for detailed notes.
Appendix E Details on the Heterogeneous Treatment Effects Across Consumers
Pre-sale Service Chatbot.
In this experiment, we focus on the comparison between consumers who are served only by the GenAI chatbot and those who are served by an auto-response indicating service unavailability. Table E1 reports the heterogeneous treatment effects across consumer groups. Our analysis reveals that the increases of both sales (Panel A) and conversion rate (Panel B) are more pronounced among newer consumers with shorter registration histories (Column 2), less active consumers with fewer login days (Column 4), and lighter consumers with less past spending (Column 6).
Table E1: Consumer HTE for Pre-sale Service Chatbot
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Sales
High Low High Low High Low
Panel A: Sales
Treat 0.268** 0.283** 0.318** 0.230** 0.194 0.314***
(0.134) (0.135) (0.156) (0.114) (0.166) (0.109)
%Change 13.7% 22.4% 15.0% 18.5% 8.6% 25.9%
Panel B: Conversion Rate
Treat 0.0134*** 0.0128*** 0.0128*** 0.0135*** 0.0134*** 0.0120***
(0.00340) (0.00368) (0.00386) (0.00321) (0.00407) (0.00310)
%Change 19.8% 26.1% 17.1% 29.6% 17.6% 25.4%
Observations 26,984 17,630 22,570 22,044 20,934 23,680
  • 1
    We classify consumers into high and low groups and capture their online shopping experience using three metrics. Consumers in the low group are considered relatively inexperienced, as determined by meeting the following pre-experiment criteria: (1) a registration duration below the platform-wide median; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
  • 2
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
  • 3
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement.
Table E2 indicates the heterogeneous treatment effect of refining search query with GenAI on different consumer groups. For both sales (Panel A) and conversion rate (Panel B), the benefits are only significant for inexperienced consumers (Columns 2, 4 and 6). Since consumers with limited online experience often struggle to effectively articulate their needs through query-based searches, they tend to benefit more from enhanced match quality achieved by utilizing GenAI to translate and refine consumer queries.
Table E2: Consumer HTE for Search Query Refinement
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Sales
High Low High Low High Low
Panel A: Sales
Treat 0.0360 0.0867** 0.0203 0.104*** -0.0216 0.106***
(0.0526) (0.0380) (0.0566) (0.0303) (0.0757) (0.0291)
%Change 1.27% 5.01% 0.63% 8.16% -0.56% 7.46%
Panel B: Conversion Rate
(0.000697) (0.000487) (0.000675) (0.000482) (0.000907) (0.000421)
%Change -0.09% 2.79% 0.43% 2.32% 0.44% 1.93%
Observation 813,317 1,036,065 890,044 959,338 596,721 1,252,661
  • 1
    Please refer to Table E1 for detailed notes.
Product Description.
In Table E3, we observe similar pattern. There is a significant and more pronounced sales increase for inexperienced consumers (Columns 2, 4 and 6). Augmenting product descriptions with GenAI-generated content substantially enhanced the sufficiency and clarity of product information, thereby motivating consumer purchase decisions, particularly among less experienced consumers.
Table E3: Consumer HTE for Product Description
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Sales
High Low High Low High Low
Panel A: Sales
Treat 0.00671 0.0130*** -0.00606 0.0263*** 0.000224 0.0206***
(0.00761) (0.00455) (0.00655) (0.00522) (0.00718) (0.00423)
%Change 1.09% 3.06% -1.02% 6.24% 0.03% 5.63%
Panel B: Conversion Rate
(0.000312) (0.000229) (0.000284) (0.000245) (0.000292) (0.000233)
%Change 1.10% 1.39% 0.65% 2.06% 1.05% 1.59%
Observation 2,035,278 2,737,659 2,322,437 2,450,500 2,386,336 2,386,601
  • 1
    Please refer to Table E1 for detailed notes.
Marketing Push Message.
Table E4 presents the buyer-level heterogeneous treatment effects of the Marketing Push Message experiment. For sales (Panel A), we find a significant increase among consumers with lower spending in the 30-day pre-experiment window (Column 6). For conversion rate (Panel B), the analysis indicates that only inexperienced consumers derive significant benefits from GenAI-driven marketing push activities (Columns 2, 4, and 6).
Table E4: Consumer HTE for Marketing Push Message
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Sales
High Low High Low High Low
Panel A: Sales
Treat 0.000269 0.00128 0.000693 0.000195 -0.00250 0.00238***
(0.00115) (0.00112) (0.00136) (0.00100) (0.00161) (0.000712)
% Change -0.9% 6.6% 2.5% 0.8% -5.9% 20.6%
Panel B: Conversion Rate
(0.0000301) (0.0000312) (0.0000345) (0.0000281) (0.0000416) (0.0000210)
% Change 1.8% 5.0% 0.7% 4.8% -2.2% 14.2%
Observations 7,959,821 5,755,707 5,796,905 7,918,623 6,085,062 7,630,466
  • 1
    Please refer to Table E1 for detailed notes.
Appendix F Details on the Heterogeneous Treatment Effects Across Products

### v3

@@ Appendix C Main Results: Model, Estimation, and Additional Outcomes
(1) (2) (3) (4) (5) (6)
Clicks Orders Sales Conversion Rate Cart Value Purchase Frequency
Treat 0.00233 0.000600** 0.0104** 0.000554*** 0.0944 -0.00186
(0.00183) (0.000278) (0.00417) (0.000187) (0.0805) (0.00320)
%Change 0.12% 1.08% 2.05% 1.27% 0.82% -0.15%
Observations 4,772,937 4,772,937 4,772,937 4,772,937 210,155 210,155
R-squared 0.055 0.008 0.002 0.008 0.010 0.021
@@
    “Clicks” stands for the number of product clicks into product detail pages. “Orders” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Purchase Frequency” refers to the number of orders per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Table C6: Product Description: Heterogeneity by Length of Original, Human-generated Description
(1) (2) (3) (4) (5) (6)
Sales Sales Conversion Rate Conversion Rate Cart Value Cart Value
$\leq$50 words >>50 words $\leq$50 words >>50 words $\leq$50 words >>50 words
Treat 0.0115** 0.0001 0.00047** 0.00044* 0.367 -0.188
(0.0050) (0.0064) (0.00022) (0.00026) (0.235) (0.198)
%Change 6.50% 0.03% 2.43% 1.60% 3.99% -1.50%
P-value 0.0536 0.5490 0.0653
Observations 1,628,409 1,628,409 1,628,409 1,628,409 31,759 45,500
R-squared 0.000 0.001 0.003 0.005 0.001 0.004
  • 1
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
  • 2
    $\leq$50 words” refers to descriptions containing fewer than 50 words, while “>>50 words” refers to descriptions containing more than 50 words.
  • 3
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. The p-values are from seemingly unrelated estimations testing the equality of % treatment effects across the two groups. *** p<<0.01, ** p<<0.05, * p<<0.1.
Marketing Push Message
@@
Table C7 shows that the use of AI-generated marketing messages leads to increases in both consumer engagement and purchases. Specifically, clicks increase by 3.1%, orders by 2.8%, and total purchase amount by 1.6% (Columns 1–3). In addition, the probability that a consumer makes a purchase rises by 3%, while cart value and purchase frequency remain statistically unchanged (Columns 4–6). Taken together, these results suggest that GenAI enhances the effectiveness of marketing content primarily by expanding consumer participation, consistent with its ability to unlock personalization at scale in settings where human-generated content is constrained by limited resources.
@@
Conversion Cart Purchase
Clicks Orders Sales Rate Value Frequency
Treat 0.000529** 0.00005* 0.000402 0.000048*** -0.206 -0.013
@@
%Change 3.1% 2.8% 1.6% 3.0% -1.32% -0.12%
Observations 13,715,528 13,715,528 13,715,528 13,715,528 22,425 22,425
R-squared 0.000 0.000 0.000 0.000 0.000 0.000
@@
    “Clicks” stands for the number of clicks on the marketing messages. “Orders” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase. “Purchase Frequency” refers to the number of orders per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@
Views Clicks Sales Conversion Rate Cart Value
@@
    “Views” represents the number of times the product advertisement is viewed on Google. “Clicks” refers to the number of times the product advertisement is clicked on Google. “Sales” represents the total expenditure on product orders. “Conversion Rate” measures users’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per product, conditional on the product being purchased.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@
Post-purchase Outcomes
In Table C9, we provide evidence on two important post-purchase measures of consumer satisfaction with the products, namely product return rates and customer ratings. We do not find evidence that GenAI applications deteriorate such outcomes.
Table C9: Average Treatment Effects of GenAI Adoption on Product Returns and Reviews
Return Rate Positive Review Rate
(1) (2) (3) (4) (5) (6)
Business Workflow Coefficient % Change Obs Coefficient % Change Obs
Pre-sale Service Chatbot
0.0208
2.08%
3,076
0.0452*
4.52%
1,008
Search Query Refinement
-0.0141
-1.4%
163,381
0.0031
0.3%
26,209
Product Description
-0.0386***
-3.9%
210,155
-0.0042
-0.4%
31,916
Marketing Push Message
-0.114***
-11.4%
22,425
-0.0174
-1.74%
2,853
  • 1
    “Return Rate” is defined as the share of orders that are returned. “Positive Review Rate” is defined as the share of rated orders that receive four- or five-star ratings on a five-point scale. The estimation of Return Rate is conditional on consumers with orders, while the estimation of Positive Review Rate is conditional on consumers provide ratings. Such data are not available for the Google Advertising Title workflow.
  • 2
    Columns (1) and (4) report the estimated coefficients, with standard errors in parentheses. Columns (2) and (5) report % Change, calculated as the treatment effect divided by the control group mean. Columns (3) and (6) report the number of observations. *** p<<0.01, ** p<<0.05, * p<<0.1.
Appendix D Details on Consumer Treatment Effect Heterogeneity
Pre-sale Service Chatbot
In this experiment, we focus on the comparison between consumers who are served only by the GenAI chatbot and those who are served by an auto-response indicating service unavailability. Table D1 reports the heterogeneous treatment effects across consumer groups. Our analysis reveals that the gains in sales are more pronounced among newer consumers with shorter registration histories (Column 2), less active consumers with fewer login days (Column 4), and consumers with less past spending (Column 6). However, differences across consumer groups are generally not statistically significant, with the exception of sales across past purchases, for which the difference is significant at the 10% level.
Table D1: Consumer HTE for Pre-sale Service Chatbot
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Purchases
High Low High Low High Low
Panel A: Sales
Treat 0.268** 0.283** 0.318** 0.230** 0.160 0.343***
(0.134) (0.135) (0.156) (0.114) (0.180) (0.105)
%Change 13.6% 22.4% 15.0% 18.5% 6.86% 27.7%
P-value 0.4568 0.7532 0.0612
Panel B: Conversion Rate
Treat 0.0134*** 0.0128*** 0.0128*** 0.0135*** 0.0129*** 0.0130***
(0.00340) (0.00368) (0.00386) (0.00321) (0.00434) (0.00300)
%Change 19.7% 26.1% 17.1% 29.6% 16.5% 27.0%
P-value 0.4579 0.1218 0.1976
Panel C: Cart Value
Treat -1.469 -0.744 -0.516 -2.331 -2.470 0.157
(1.275) (1.770) (1.387) (1.525) (1.568) (1.320)
% Change -5.07% -2.88% -1.82% -8.56% -8.27% 0.61%
P-value 0.7787 0.3702 0.2710
Observation (Panel A&B) 26,984 17,630 22,570 22,044 18,608 26,006
Observation (Panel C) 2,066 1,010 1,877 1,199 1,609 1,467
  • 1
    We classify consumers into high and low groups and capture their online shopping experience based on registration duration, 30-day pre-experiment login days and 30-day pre-experiment purchase. For each indicator, consumers in the low group are defined as relatively inexperienced if they fall below the median of the corresponding distribution.
  • 2
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
  • 3
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement
Table D2 reports the heterogeneous treatment effect of refining search query with GenAI on different consumer groups. For sales, the benefits are significant and larger for inexperienced consumers (Columns 2, 4, and 6), while the effects for more experienced consumers are smaller and insignificant (Columns 1, 3, and 5). The differences between these two consumer groups are statistically significant when consumers are classified by past purchases and past login days. When focusing on conversion rates, we also find statistically significant differences between consumer groups defined by years since registration. Since consumers with limited online experience often struggle to effectively articulate their needs through query-based searches, they tend to benefit more from enhanced match quality achieved by applying GenAI to the semantic translation consumer queries.
Table D2: Consumer HTE for Search Query Refinement
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Purchases
High Low High Low High Low
Panel A: Sales
Treat 0.0360 0.0867** 0.0203 0.104*** -0.0216 0.106***
(0.0526) (0.0380) (0.0566) (0.0303) (0.0757) (0.0291)
%Change 1.27% 5.01% 0.63% 8.16% -0.56% 7.46%
P-value 0.2137 0.0036 0.0047
Panel B: Conversion Rate
(0.000697) (0.000487) (0.000675) (0.000482) (0.000907) (0.000421)
%Change -0.09% 2.79% 0.43% 2.32% 0.44% 1.93%
P-value 0.0023 0.0421 0.1122
Panel C: Cart Value
Treat 0.329 0.440 -0.015 1.126** -0.355 1.224***
(0.431) (0.526) (0.452) (0.460) (0.484) (0.450)
%Change 1.34% 1.69% -0.05% 5.36% -1.35% 5.13%
P-value 0.8965 0.0363 0.0139
Observation (Panel A&B) 813,317 1,036,065 890,044 959,338 596,721 1,252,661
Observation (Panel C) 93,362 70,019 104,531 58,850 88,135 75,246
  • 1
    Refer to Table D1 for detailed notes.
Product Description
In Table D3, we observe a similar pattern to search. There is a significant and more pronounced sales increase for inexperienced consumers (Columns 2, 4, and 6). The differences between high- and low-group consumers are statistically significant for consumers classified by past purchases and past login days. Augmenting product descriptions with AI-generated content substantially enhances the sufficiency and clarity of product information, thereby motivating consumer purchase decisions, particularly among less experienced consumers.
Table D3: Consumer HTE for Product Description
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Purchases
High Low High Low High Low
Panel A: Sales
Treat 0.00671 0.0130*** -0.00606 0.0263*** 0.000224 0.0206***
(0.00761) (0.00455) (0.00655) (0.00522) (0.00718) (0.00423)
%Change 1.09% 3.06% -1.02% 6.24% 0.03% 5.63%
P-value 0.2525 0.0001 0.0005
Panel B: Conversion Rate
(0.000312) (0.000229) (0.000284) (0.000245) (0.000292) (0.000233)
%Change 1.10% 1.39% 0.65% 2.06% 1.05% 1.59%
P-value 0.7441 0.1194 0.5235
Panel C: Cart Value
Treat 0.003 0.184* -0.180 0.438*** -0.127 0.447***
(0.126) (0.098) (0.112) (0.115) (0.115) (0.098)
%Change 0.03% 1.62% -1.50% 3.96% -1.06% 4.09%
P-value 0.2496 0.0001 0.0001
Observation (Panel A&B) 2,035,278 2,737,659 2,322,437 2,450,500 2,386,336 2,386,601
Observation (Panel C) 107,098 103,057 115,708 94,447 129,462 80,693
  • 1
    Refer to Table D1 for detailed notes.
Marketing Push Message
In Table D3, we observe a pattern similar to that in Search Query Refinement. Sales increases are larger and statistically significant for less experienced consumers (Columns 2, 4, and 6). Differences between high- and low-experience consumers are statistically significant when consumers are classified by past purchases and past login days. Augmenting product descriptions with AI-generated content improves the sufficiency and clarity of product information, thereby increasing the likelihood of purchase, particularly among less experienced consumers.
Table D4: Consumer HTE for Marketing Push Message
(1) (2) (3) (4) (5) (6)
Registered Years Past Login Days Past Purchases
High Low High Low High Low
Panel A: Sales
Treat -0.000269 0.00128 0.000693 0.000195 -0.00250 0.00238***
(0.00115) (0.00112) (0.00136) (0.00100) (0.00161) (0.000712)
% Change -0.91% 6.65% 2.47% 0.84% -5.89% 20.6%
P-value 0.4519 0.8606 0.0003
Panel B: Conversion Rate
(0.0000301) (0.0000312) (0.0000345) (0.0000281) (0.0000416) (0.0000210)
% Change 1.76% 4.97% 0.71% 4.82% -2.22% 14.2%
P-value 0.4373 0.2809 0.0000
Panel C: Cart Value
Treat -0.432 0.225 0.285 -0.573 -0.6 0.828
(0.578) (0.731) (0.714) (0.583) (0.558) (0.761)
% Change -2.62% 1.60% 1.75% -3.8% -3.76% 5.62%
P-value 0.4956 0.3551 0.1286
Observations (Panel A&B) 7,959,821 5,755,707 5,796,905 7,918,623 6,085,062 7,630,466
Observations (Panel C) 14,351 8,074 10,026 12,399 16,025 6,400
  • 1
    Refer to Table D1 for detailed notes.
Appendix E Details on Seller Treatment Effect Heterogeneity
Search Query Refinement
@@ Appendix E Details on Seller Treatment Effect Heterogeneity
Table E1 reports the heterogeneous impact on high- versus low-type of sellers classified based on three metrics: annual past sales, tenure on the platform, and number of sub-accounts. Small sellers with lower transaction volumes (Column 2), shorter operational histories (Column 4), and fewer sub-accounts (Column 6) experience a significant and larger increase in sales from treated consumers. By contrast, sales on larger, tenured, and scaled sellers show no significant change. The differences between these two seller groups are statistically significant when sellers are divided by operation years and the number of sub-accounts. When using conversion rates, we also find statistically significant differences between seller groups defined by annual sales. Thus, the GenAI-powered search query refinement generates more values among small sellers.
Table E1: Seller HTE for Search Query Refinement
@@
%Change 0.21% 1.69% 0.76% 1.47% 0.49% 1.85%
P-value 0.0573 0.3493 0.0738
Panel C: Cart Value
Treat 0.550 0.238 -0.042 0.773 -0.054 0.613
(0.554) (0.267) (0.287) (0.486) (0.325) (0.426)
%Change 2.06% 1.34% -0.22% 3.29% -0.26% 2.75%
P-value 0.7711 0.1484 0.2069
Observation (Panel A&B) 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382
Observation (Panel C) 76,536 115,804 108,790 85,195 91,774 100,643
@@
    We classify sellers into high and low groups based on three different proxies. Low-group sellers are generally small sellers, as defined by meeting the following pre-experiment criteria: (1) accounting for the bottom 50% cumulative share of total sales; (2) having operated on the platform for fewer than five years; or (3) maintaining fewer than three sub-accounts for their online store.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Marketing Push Message
In Table E2, the differences between high- and low-group sellers are all statistically insignificant.
Table E2: Seller HTE for Marketing Push Message
@@
% Change 3.27% 2.69% 2.39% 3.97% -0.09% 4.20%
P-value 0.8299 0.1826 0.1485
Panel C: Cart Value
Treat -0.221 -0.195 0.430 -1.403 1.237 -0.829
(0.734) (0.564) (0.485) (0.918) (0.835) (0.540)
% Change -1.38% -1.27% 3.22% -7.1% 10.1% -4.87%
P-value 0.9852 0.0817 0.0467
Observation (Panel A&B) 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528
Observation (Panel C) 10,212 12,213 14,405 8,020 6,410 16,015
@@
    Refer to Table E1 for detailed notes.
Google Advertising Titles
Table E3 shows that the differences between high- and low-group sellers are all statistically insignificant.
Table E3: Seller HTE for Google Advertising Titles
@@
% Change -6.33% -0.33% -2.25% -3.73% -4.61% -3.18%
P-value 0.3549 0.8248 0.8823
Panel C: Cart Value
Treat -0.0528 -1.588 -0.990 -0.697 -0.960 -0.749
(1.372) (1.425) (1.896) (1.162) (3.811) (1.016)
% Change -0.16% -4.41% -2.84% -2.09% -2.28% -2.28%
P-value 0.4433 0.8971 0.9997
Observations (Panel A&B) 622,083 621,933 397,339 846,677 140,005 1,104,011
Observation (Panel C) 2,343 2,468 1,411 3,400 468 4,343
@@
    Refer to Table E1 for detailed notes.
Appendix F Details on Product Treatment Effect Heterogeneity